|[Date Prev][Date Next][Thread Prev][Thread Next][Date Index][Thread Index]
Cassini Significant Events 02/23/11 - 03/01/11
The most recent spacecraft telemetry was acquired on Mar. 1 from the Deep Space Network tracking complex at Goldstone, California. The Cassini spacecraft is in an excellent state of health and all subsystems are operating normally. Information on the present position and speed of the Cassini spacecraft may be found on the "Present Position" page at: http://saturn.jpl.nasa.gov/mission/presentposition/.
Wednesday, Feb. 23 (DOY 054)
This week in science, Cassini was outbound from Saturn periapsis in a sequence segment focused on Saturn measurements. The Composite Infrared Spectrometer (CIRS) performed a mid-infrared mapping observation, which helps determine Saturn’s upper troposphere and tropopause temperatures at a spatial resolution of about two degrees of latitude and longitude on the planet. CIRS also made several observations to measure oxygen compounds in Saturn’s stratosphere as a function of latitude, while the Visual and Infrared Mapping Spectrometer (VIMS) rode along. Imaging Science (ISS) and VIMS performed a series of Saturn cloud-tracking observations. The Optical Remote Sensing (ORS) teams took another two looks at Titan as part of their cloud monitoring campaign. Additional science observations included an Ultraviolet Imaging Spectrograph Saturn extreme ultraviolet/far-ultraviolet observation involving slow scans across the visible hemisphere to form spectral images. The Magnetospheric and Plasma Science (MAPS) instruments were occupied with a magnetospheric survey campaign, and ending the week the Radio Science (RSS) performed a one-hour boresight calibration.
Thursday, Feb. 24 (DOY 055)
The final sequence approval meeting for S67 was held today. The S67 background sequence goes active on March 7.
Five Instrument Expanded Block (IEB) files were uplinked to the spacecraft today in support of S67. After reviewing the memory readouts, the uplink team verified that the files have been properly loaded into the solid-state recorder.
Friday, Feb. 25 (DOY 056)
The Reaction Wheel Assembly (RWA)-4 to 3 transition activities are still in work. RWA-3 will be brought on-line in place of RWA-4 on March 7 in preparation for S67.
Saturday, Feb. 26 (DOY 057)
The Downlink Ground System (DGS) team was provided with an acceptable gplot workaround which is already part of operations. The gplot software is a Multi-Mission Ground System and Services tool that plots telemetry channel values.
The DGS Team successfully completed setup of the Science Operations Planning Computer (SOPC) client for testing of the Distributed Object Manager on the Network File System.
Sunday, Feb. 27 (DOY 058)
The VIMS L1A data product generation software encountered problems creating a VIMS data cube. Instrument Operations/ Multi-Mission Image Processing Laboratory (IO/MIPL) personnel manually bypassed the problematic data, and are investigating the cause of the problem.
IO/MIPL loaded predicts for S67 into the MIPL Cassini database. This primes the automated system to expect the downlinks and instrument products.
Monday, Feb. 28 (DOY 059)
Port 1 products were due today as part of the S69 Sequence Implementation Process (SIP). The products will be merged and sent out to the flight team for review.
The Science Forum for S69 was held today. Topics included an overview of science planned for this sequence followed by highlights, unique activities, and highest priority observations as described by the Target Working Team (TWT) and Orbiter Science Team (OST) leads, with comments from the Investigation Scientists and other instrument team representatives.
Tuesday, March 1 (DOY 060)
There was an increase in the Radio Frequency Subsystem (RFS) Static Phase Error (SPE) today - the Deep Space Transponder A rose 3 kHz during the Titan 74 (T-74) flyby last week. The Spacecraft Operations team requested the DSN to raise the best lock frequency by 3 kHz. This change will be implemented by the DSN on March 3, during the backup OTM pass on DOY 062.
The maneuver approval meeting for Orbit Trim Maneuver (OTM) #279 was held today. OTM-279 is scheduled to execute tomorrow morning using the Reaction Control Subsystem thrusters. To save delta-V, the Navigation team intends to re-target the T-75 aimpoint by about 13 km. The science teams have evaluated a trajectory implementing the new aimpoint and confirmed their observations will not be impacted.
To unsubscribe from Cassini Spacecraft Updates or to subscribe with a different email address, visit here:
This email address is not monitored. For comments and questions, please contact Cassini Outreach here:
Visit the JPL Cassini home page for more information about the Cassini Project:
[NASA News] [JPL Home] [JPL News] [NASA KSC] [NTSB] [Hot Springs] [Yosemite News] [NSF] [NASA News] [Bake Sale for NASA] [Telescopes]